OpenStudy (anonymous):

1. What number would you have to add to both sides to complete the Square? X^2-16x=-8 A.-16 B.16 C.64 D.-64 2. Solve by using the perfect squares method X^2-12x+36=0 A. -18 B. 18 C. -6 D. 6

OpenStudy (welshfella):

divide the coefficiebt of x by 2 . In this cast the coefficient is -16 than write it as ( x - a)^2 - c so that this equals x^2 - 16x

OpenStudy (anonymous):

So #1 is B!?

OpenStudy (welshfella):

no what is -16 / 2 ?

OpenStudy (anonymous):

-8

OpenStudy (welshfella):

- this will give yiu the value of a in (x + a)^2

OpenStudy (welshfella):

right so we have (x - 8)^2 + b now (x - 8)^2 = x^2 - 16x + 64 so get the left side = x^2 - 16x we have to add - 64

OpenStudy (welshfella):

and we have to add -64 to both sides

OpenStudy (anonymous):

Ok

OpenStudy (welshfella):

for number 2 if you factor the trinomial you'll find its a perfect square so you'll get ( x - a)^2 = 0

OpenStudy (welshfella):

and the root will be x = a

OpenStudy (welshfella):

youll need to use one of the square roots of 36

OpenStudy (anonymous):

6

OpenStudy (welshfella):

thats one of the square roots - but you need the other one because the middle term is -12x

OpenStudy (anonymous):

Oh okay so -6 then?

OpenStudy (welshfella):

yes so it factors to (x - 6)^2 = 0 so what is the value of x?

OpenStudy (welshfella):

x - 6 must be equal to 0, right?

OpenStudy (anonymous):

Yeah

OpenStudy (welshfella):

so what does x =?

OpenStudy (welshfella):

x - 6 = 0 so x = ? what have you got to take 6 from to get 0?

OpenStudy (anonymous):

6

OpenStudy (welshfella):

exactly so thats the answer x = 6

OpenStudy (anonymous):

Thanks so much! I'm super bad at math, so teaching just require Patience lol. Thanks. So #1 is C and #2 is d. Correct?

OpenStudy (anonymous):

Must*

OpenStudy (anonymous):

It was correct. Thanks